Can [ i ] be a column of a unitary matrix?

[-i ]

Justify your answer



1
Expert's answer
2021-07-30T09:05:33-0400

Yes, [i] can be a column at a unitary matrix as if we choose "1 \times 1" matrix. I.e i=j, meaning that row = column.

So in this case we can say that [i] as a column of a unitary matrix.


Consequently, yes, [-i] can be a column at a unitary matrix as if we choose "1 \times 1" matrix. I.e -i=-j, meaning that row = column.

So in this case we can say that [-i] as a column of a unitary matrix.